OpenStudy (anonymous):

Factor the trinomial... s^2-2sf-63f^2

jhonyy9 (jhonyy9):

-2=-9+7 -63=-9*7

OpenStudy (anonymous):

so what is the answer i am confused sorry

jhonyy9 (jhonyy9):

(s-9f)(s+7f)

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

ok